# Improving Generalization in Game Agents with Data Augmentation in Imitation Learning

### 21 Dec 2023 v2

@@ V-B Evaluation
Figure 5: Relative performance, averaged over all environments, is consistently poor or failing for the bottom group of models. Abbreviations: gauss (Gaussian noise), uni (uniform noise) sca (scaling), sm (state-mixup), drc (continuous dropout), drs (semantic dropout), rX (X percentage of the original dataset used), and eX ($\sigma=3\cdot 10^{-\text{X}}$). Additional details about the augmentations are provided in Section III.
@@ VI Conclusion and discussions
In this paper we have conducted a comprehensive study of data augmentation in imitation learning. Training self-learning agents with either reinforcement learning or imitation learning, involves a generalization–efficiency trade-off: we can increase generalization, but at the cost of sample efficiency. Data augmentation can help improve generalization while maintaining sample efficiency. Building on the success of data augmentation in supervised learning and reinforcement learning, we investigated data augmentation techniques for feature-based imitation learning and their effects on the training of imitative agents. We evaluated our agents in $4$ distinct 3D test environments. The main findings of this paper are twofold. First, data augmentations can indeed improve performance in an imitation learning setting. The imitative agents demonstrated improved generalization in previously unseen situations when trained on an augmented dataset. Second, our study indicates that certain combinations of data augmentations consistently enhance performance across a variety of generalization environments, even if the performance of individual augmentations may vary depending on the parameters used.

### 07 Apr 2024 v3

@@ V-B Evaluation
Figure 5: Relative performance, averaged over all environments, is consistently poor or failing for the bottom group of models. Abbreviations: gauss (Gaussian noise), uni (uniform noise) sca (scaling), sm (state-mixup), drc (continuous dropout), drs (semantic dropout), rX (X percentage of the original dataset used), and eX ($\sigma=3\cdot 10^{-\text{X}}$).
@@ VI Conclusion and discussions
In this paper we have conducted a comprehensive study of data augmentation in imitation learning. Training self-learning agents with either reinforcement learning or imitation learning, involves a generalization–efficiency trade-off: we can increase generalization, but at the cost of sample efficiency. Data augmentation can help improve generalization while maintaining sample efficiency. Building on the success of data augmentation in supervised learning and reinforcement learning, we investigated data augmentation techniques for feature-based imitation learning and their effects on the training of imitative agents. We evaluated our agents in four distinct 3D test environments. The main findings of this paper are twofold. First, data augmentations can indeed improve performance in an imitation learning setting. The imitative agents demonstrated improved generalization in previously unseen situations when trained on an augmented dataset. Second, our study indicates that certain combinations of data augmentations consistently enhance performance across a variety of generalization environments, even if the performance of individual augmentations may vary depending on the parameters used.